# Correcting Annotator Bias in Training Data: Population-Aligned Instance Replication (PAIR)

### v2

@@ Appendix C Ablation Study 1: HS Results
Figure 4: Percentage of tweets annotated as HS, by dataset and bias ($\beta$)
HS is rare in our dataset (16.7% of HS labels are yes, Table 1). And, our simulation strategy overrepresents $A$ annotators in the two Non-representative data sets, who are less likely to perceive HS (Table 2). For these reasons, a higher proportion of $p_{i,HS}^{A}$ are 00 while the $p_{i,HS}^{B}$ probabilities continue to increase. This issue leads the proportion of “yes” labels in the Representative and Adjusted datasets to increase with $\beta$ in the HS data set, which have more B labels than the unadjusted datasets.
Results.
Figure 5 (1) contains the ACB results and Figure 5 (2) the F1 score results for the HS datasets. The PAIR approach does not improve calibration or accuracy: the adjusted model performs similarly to the Non-representative models. This effect is likely due to the combination of label rarity and our simulation design. With few positive labels to begin with, the impact of varying annotator characteristics through our $\beta$ parameter may be overwhelmed by the baseline scarcity of hate speech annotations. Ablation Study 2 (Appendix §D) addresses this point.
Figure 5: Metrics for HS models, by dataset and bias ($\beta$)
Appendix D Ablation Study 2: Difficult Tweets
Our simulations assumed that all tweets are impacted the same way (Eq. (2)), which is an oversimplification. More realistically, annotator characteristics likely have more impact for ambiguous tweets. For this reason, we repeat model training and recompute metrics for those tweets where $0.4\leq p_{i}\leq 0.6$. This approach not only focuses on those tweets where annotator characteristics likely play a larger role, it also eliminates the floor and ceiling effects in Eq. 2). The filtered datasets contain 267 (OL) and 360 (HS) tweets.
Figures 6 and 7 show the results for two metrics (ACB, F1) for filtered OL and HS labels. In Figure 6(1): the Representative and Adjusted models have similar ACB and are lower than the Non-representative models. The F1 scores do not show differences between the models. These results are similar to those on the full set of tweets (Figure 3). In the two HS figures (7), we see signs that the Representative and Adjusted models perform similarly, and better than the two Non-representative models, on both metrics. These results are more promising than those on the full set of tweets (Figure 5).
@@ Appendix D Ablation Study 2: Difficult Tweets
Figure 6: Metrics for OL models for filtered tweets ($0.4\leq p_{i,OL}\leq 0.6$), by dataset and bias ($\beta$)
@@
Figure 7: Metrics for HS models for filtered tweets ($0.4\leq p_{i,HS}\leq 0.6$), by dataset and bias ($\beta$)

### v3

@@ 5.1 OL Models
Figure 9: Percentage of OL instances on difficult tweets ($0.4\leq p_{i,OL}\leq 0.6$) by dataset and bias ($\beta$)
@@
Figure 10: Percentage of HS instances on difficult tweets ($0.4\leq p_{i,HS}\leq 0.6$) by dataset and bias ($\beta$)
Accuracy.
Figure 6 compares the models’ F1 scores. In contrast to Figure 5, we do not see strong differences between the models trained on the different datasets. For all datasets, model performance declines with $\beta$: as the amount of bias in the annotations increases, the models are less able to predict the binary OL label.
Because the F1 metric focuses on binary predictions, it is less sensitive to training biases than calibration metrics like ACB, which more explicitly capture biases through prediction scores. In decision-making, miscalibrated predictions can have harmful consequences when, for example, hateful content remains undetected Van Calster et al. (2019). These findings suggest that calibration metrics provide a clearer view of the impact of annotators on models: binary classification metrics can obscure such effects.
5.2 HS Models
Figure 7 contains the ACB results and Figure 8 the F1 score results for the HS models trained on each dataset. Though the adjusted model roughly tracks the representative models for ACB, there is instability in the results. All models show lower average ACB values than the representative model across a wide range of the bias offset (0.10 - 0.20). The PAIR approach does not improve calibration or accuracy: the adjusted model performs similarly to the Non-representative models. This effect is likely due to the combination of label rarity and our simulation design. With few positive annotations to begin with, the impact of the $\beta$ parameter and the overrepresentation of the A annotators may be overwhelmed by the baseline scarcity of hate speech annotations. Calibration metrics can be less reliable with rare classes Zhong et al. (2021).
5.3 Sensitivity Analysis: Difficult Tweets
Our simulations assumed that all annotator type impacts all tweets the same way (Eq. 5), which is an oversimplification. More likely, annotator characteristics have more impact for ambiguous tweets. For example, prior research in the psychology literature on judgment under uncertainty suggests that people draw more heavily on personal heuristics when interpreting unclear or underspecified information Tversky and Kahneman (1974). For this reason, we repeat model training and recompute metrics for those tweets where $0.4\leq p_{i}\leq 0.6$. Subsetting the tweets in this way also eliminates the floor and ceiling effects in Eq. 5. The filtered datasets contain 267 (OL) and 360 (HS) tweets. The proportions of OL and HS annotations are stable for the Representative and Adjusted sets and decrease for the Non-representative sets as we increase the bias offset (Figures 9 and 10). This mimics the trend for OL in the full dataset (Figure 3).
@@ 5.3 Sensitivity Analysis: Difficult Tweets
Figure 11: ACB scores for OL Models, on difficult tweets ($0.4\leq p_{i,OL}\leq 0.6$), by dataset and bias ($\beta$)
Figure 12: F1 scores for OL Models, on difficult tweets ($0.4\leq p_{i,OL}\leq 0.6$), by dataset and bias ($\beta$)
Figures 11, 12, 13, and 14 show results for two metrics (ACB, F1) for filtered OL and HS annotations. In Figure 11, the Representative and Adjusted models have similar ACB and are lower than the Non-representative models. The F1 scores do not show differences between the models. These results are similar to those on the full set of tweets (Figures 5 and 6). In the two HS figures (13, 14), we see signs that the Representative and Adjusted models perform similarly, and better than the two Non-representative models, on both metrics. These results are more promising than those on the full set of tweets (Figures 7 and 8) and support our hypothesis that the rarity of HS annotations contributed to the lack of positive results for the PAIR approach in §5.2. The PAIR algorithm works well with difficult tweets, which is where it is likely most needed.
Figure 13: ACB scores for HS Models, on difficult tweets ($0.4\leq p_{i,HS}\leq 0.6$), by dataset and bias ($\beta$)
Figure 14: F1 scores for HS Models, on difficult tweets ($0.4\leq p_{i,HS}\leq 0.6$), by dataset and bias ($\beta$)
6 Discussion & Recommendations
Our experimental results show the OL prediction models perform less well when trained on data from non-representative annotator pools (RQ1), and simple statistical adjustments can improve model calibration without collecting additional annotations or involving additional annotators (RQ2). While PAIR’s impact was harder to assess for the rare HS class, PAIR did improve calibration of both the OL and HS models when trained on difficult tweets. These findings establish a promising bridge between survey statistics and machine learning – offering a practical approach to make AI systems more representative of and responsive to the populations they serve, particularly for tasks involving subjective human judgments.
We recommend the following four steps to reduce bias due to non-representative annotator pools:
  1. 1)
  2. 2)
    Collect these characteristics from annotators and gather corresponding population-level data from national censuses or high-quality surveys. [^5] Collection and release of annotator characteristics or weights derived from them may raise confidentiality concerns. The survey literature offers advice for sharing sensitive data (see Karr, 2016, for a review). Collecting annotator characteristics may also require involvement of Institutional Review Boards or other participant protection organizations Kaushik et al. (2024).
  3. 3)
  4. 4)
    Use these weights in model training. Our simple replication approach showed promise, and future work should test more sophisticated weighting approaches.
7 Limitations
Our study explores bias-aware data simulation and evaluation in a controlled setting, which necessarily involves simplifying assumptions and methodological constraints. We outline key areas where future work could broaden the applicability and robustness of our findings.
Stylized Biases and Simulated Data.
Our simulation makes strong assumptions about annotator behavior: there are only two types of annotators, and, within each type, annotators behave similarly. Real-world annotator biases may be more nuanced or context-dependent. The simulated annotators might not be representative of a stable opinion group Mokhberian et al. (2024); Vitsakis et al. (2024). Future work could incorporate more realistic biases and refine the proposed simulations and statistical techniques.
Sampling Variability.
We have created only one version of the four datasets for each annotation type and value of $\beta$, each of which contains random draws from the Bernoulli distribution. A more traditional statistical approach would create multiple versions of the datasets and train models on each one, to average over the sampling variability. Though limited by computational constraints in this work, future work could take on a more expansive simulation. As discussed, we used five seeds in model training.
Need for Population Benchmarks and Annotator Characteristics.
PAIR requires high quality benchmark information about the relevant population. These benchmarks might come from national statistical offices or national surveys. Annotators must provide accurate data on the same characteristics available in the benchmark data. Unfortunately, annotators sometimes do not provide accurate information Chandler and Paolacci (2017); Huang et al. (2023). In addition, theory demonstrates that bias will be reduced only when the characteristics used in weighting correlate with the annotations (Eckman et al., 2024). In our simulation, differences in annotations were driven solely by group membership ($A$, $B$). In the real world, it is challenging to know what characteristics impact annotation behavior for a given task and to find good benchmarks for those characteristics.
Generalization Beyond Task Types.
Evaluation Metrics.
8 Ethical Considerations
In this simulation study, we experiment on a publicly available dataset collected in our previous study Kern et al. (2023), which contains offensive and hateful tweets. We do not support the views expressed in these tweets. The simulation study itself does not collect any new data or raise any ethical considerations.
Acknowledgments